# Formalizing zeta and L-functions in Lean

### v3

@@ Abstract.
D.L. gratefully acknowledges the support of the European Research Council through the Horizon 2020 Excellent Science programme (Consolidator Grant “ShimBSD: Shimura varieties and the BSD conjecture”, grant ID 101001051). M.S. thanks the Institute for Mathematical Research (FIM) at ETH Zürich for enabling him to spend a week at ETH
1. Introduction
1.1. Mathematical background
@@
1.2. Results formalized in this project
@@ 1.3. Related work
2. Implementing $L$-series in Mathlib
@@
It turned out that it was not easily possible to set up a general coercion from ArithmeticFunction $R$ to ArithmeticFunction $R^{\prime}$ when there is a coercion from $R$ to $R^{\prime}$, because the necessary coercion instance would have too many free parameters; and hard-coding the relevant cases $\mathbb{N}\hookrightarrow\mathbb{Z}\hookrightarrow\mathbb{R}\hookrightarrow\mathbb{C}$ led to a considerable amount of code duplication. See the discussion on the Lean Zulip chat [^2] https://leanprover.zulipchat.com/#narrow/channel/144837-PR-reviews/topic/.2310725.20and.20.2310728.3A.20L-series/near/422590170 for details on this and also on how we arrived at the current design, which is described below.
@@
3. From Fourier analysis to zeta functions
3.1. Fourier analysis
@@
3.2. Jacobi’s theta function
@@ 3.3. Generalizations
4. Special values of $L$-functions
@@
This approach generalizes immediately to computing values of Hurwitz zeta functions. The result is most tidily stated in terms of values at negative integers (HurwitzZeta.hurwitzZeta_neg_nat):
@@
5. Pitfalls: totalizing functions
Mathematically, the natural domain of $\zeta(s)$ is the “punctured complex plane” $\mathbb{C}-\{1\}$; there is no sensible definition of $\zeta(1)$ as an element of $\mathbb{C}$. However, in formalizing proofs it is inconvenient to work with functions defined on subtypes; so the usual practice is to extend partially-defined functions to total functions by assigning junk values at the bad points, such as the convention that $1/0=0$ used by many theorem provers (including Coq, Isabelle and Lean) [Buz20].
@@ 5. Pitfalls: totalizing functions
6. Applications to primes in arithmetic progressions
@@
To prove Dirichlet’s Theorem, we consider the $L$-series associated to the function that sends $n$ to $\Lambda(n)$ (where $\Lambda$ is the von Mangoldt function) when $n\equiv a\bmod q$ and to zero otherwise. This $L$-series is a linear combination of logarithmic derivatives of $L$-series associated to the Dirichlet characters mod $q$: ArithmeticFunction.vonMangoldt.LSeries_residueClass_eq. This step in the proof uses the orthogonality relations for (Dirichlet) characters, which we also contributed to Mathlib. The non-vanishing result above then implies that these logarithmic derivatives extend continuously to $\Re(s)\geqslant 1$, except for a simple pole at $s=1$ when $a$ and $q$ are coprime. The existence of this simple pole then fairly quickly leads to a proof that $\sum_{n\equiv a\bmod q}\Lambda(n)/n$ diverges, which in turn easily implies the existence of infinitely many prime numbers $p\equiv a\bmod q$. We state this final result in two versions, Nat.setOf_prime_and_eq_mod_infinite (the set of these primes is infinite) and Nat.forall_exists_prime_gt_and_eq_mod (for every natural number $n$, there exists a larger prime $p\equiv a\bmod q$).

### v4

@@ Abstract
keywords:
Riemann zeta, analytic number theory, formalization of proofs
[…]
1 Introduction
1.1 Mathematical background
@@
1.2 Results formalized in this project
@@ 1.3 Related work
2 Implementing $L$-series in Mathlib
@@
We found it was not easily possible to set up a coercion from ArithmeticFunction $R$ to ArithmeticFunction $R^{\prime}$ whenever there is a coercion from $R$ to $R^{\prime}$, because the necessary coercion instance would have too many free parameters; and hard-coding the relevant cases $\mathbb{N}\hookrightarrow\mathbb{Z}\hookrightarrow\mathbb{R}\hookrightarrow\mathbb{C}$ led to a considerable amount of code duplication. See the discussion on the Lean Zulip chat [^2] https://leanprover.zulipchat.com/#narrow/channel/144837-PR-reviews/topic/.2310725.20and.20.2310728.3A.20L-series/near/422590170 for details on this and also on how we arrived at the current design, which is described below.
@@
3 From Fourier analysis to zeta functions
3.1 Fourier analysis
@@
3.2 Jacobi’s theta function
@@ 3.3 Generalizations
4 Special values of $L$-functions
@@
This approach generalizes immediately to computing values of Hurwitz zeta functions. The result is most tidily stated in terms of values at negative integers:
@@
5 Pitfalls: totalizing functions
Mathematically, the natural domain of $\zeta(s)$ is the “punctured complex plane” $\mathbb{C}-\{1\}$; there is no sensible definition of $\zeta(1)$ as an element of $\mathbb{C}$. However, in formalizing proofs it is inconvenient to work with functions defined on subtypes; so the usual practice is to extend partially-defined functions to total functions by assigning junk values at the bad points, such as the convention that $1/0=0$ used by many theorem provers (including Coq, Isabelle and Lean) [div0].
@@ 5 Pitfalls: totalizing functions
6 Applications to primes in arithmetic progressions
@@
To prove Dirichlet’s Theorem, we consider the $L$-series associated to the function that sends $n$ to $\Lambda(n)$ (where $\Lambda$ is the von Mangoldt function) when $n\equiv a\bmod q$ and to zero otherwise. We show in ArithmeticFunction.vonMangoldt.LSeries_residueClass_eq that this $L$-series is a linear combination of logarithmic derivatives of $L$-series associated to the Dirichlet characters mod $q$. This step in the proof uses the orthogonality relations for (Dirichlet) characters, which we also contributed to Mathlib. The non-vanishing result above then implies that these logarithmic derivatives extend continuously to $\Re(s)\geqslant 1$, except for a simple pole at $s=1$ when $a$ and $q$ are coprime. The existence of this simple pole then fairly quickly leads to a proof that $\sum_{n\equiv a\bmod q}\Lambda(n)/n$ diverges, which in turn easily implies the existence of infinitely many prime numbers $p\equiv a\bmod q$. We state this final result in two versions, Nat.setOf_prime_and_eq_mod_infinite (the set of these primes is infinite) and Nat.forall_exists_prime_gt_and_eq_mod (for every natural number $n$, there exists a larger prime $p\equiv a\bmod q$).